OpenStudy (anonymous):

There is a sequence {x_n} of non-square numbers (example: 2,3,5,6,7,8,10,...). If m^2<=x_n<=(m+1)^2, then prove that m=[n+0.5], where [] is the greatest integer function.

OpenStudy (anonymous):

Can we proceed by induction here. I used induction on m and n. (clearly m and n are positive integers) the proof goes as: Base step: for n=1, 1<2<4 u get m=1 Induction hypothesis: assume p=[q+0.5] to be true. add 1 on both sides and u get p+1=[q+0.5]+1=[g+1+0.5]. But this implies that the equation to be proved holds true for p+1 when it holds true for p. So it holds true for all n, using principle of mathematical induction. Is this correct? thnx. :)

OpenStudy (anonymous):

is there any other method than induction... it will be gr8 if u come with one.. :) also please correct me if i am wrong anywhere! :)

OpenStudy (nowhereman):

I don't think this is true, because there is no given relation between x_n and n other than that the sequence contains no squares...

OpenStudy (anonymous):

yup, nowhereman thats the place where this question (and the proof) goes weird...there is no place where i m using any thing about squares!

OpenStudy (nowhereman):

Already your induction start won't work. You can not choose an arbitrary m. The proposition is that for any m, if the inequality holds, then m=[n+0.5], but you don't know what \(x_1\) is! So [n+0.5] will always be 1 but m could be any number.

OpenStudy (anonymous):

can u tell me what is n exactly here?

OpenStudy (anonymous):

if it is an integer then [n+.5] has no meaning...

OpenStudy (anonymous):

k...nowhereman..u may be correct.. i am wrong in the induction hypothesis part... but we can have that the equation holds true for all such q i.e. n (its easy to prove and on similar lines just like the one i gave)

OpenStudy (anonymous):

then the proof seems correct? (but nevertheless weird since no use of the fact that they are all non-squares)

OpenStudy (nowhereman):

No, the proposition is false. You can easily construct a sequence of non-square numbers where \(m^2 \leq x_n \leq (m+1)^2\) but \(m \neq n = [n+0.5]\).

OpenStudy (anonymous):

so how will u prove?

OpenStudy (nowhereman):

You can not prove something which is wrong...

OpenStudy (anonymous):

the question is right! how do u say its wrong?

OpenStudy (nowhereman):

take the sequence \(100n^2+1\) It certainly contains no squares. But the first element is \(101\) So for \(m=10\) you have \(m^2 = 100 \leq 101 = x_1 \leq 121 = (m+1)^2\) but you also have \(m = 10 \neq 1 = [1 + 0.5] = [n+0.5]\).

OpenStudy (anonymous):

u must have the sequence starting with 2, i.e. x_1=2,x_2=3 and so on...not from any arbitrary non-square number....

OpenStudy (anonymous):

u misinterpreted the question, nowhereman....

OpenStudy (anonymous):

or maybe i should have explicitly told...sorry :)

OpenStudy (anonymous):

u have the sequence {x_n} with x_1=2, x_2=3, x_3=5, ...

OpenStudy (anonymous):

now having known this, is my proof correct?

OpenStudy (nowhereman):

Wait, so should the proposition be "Let \((x_n)_{n=1}^∞\) be the ascending sequence of all non-square integers."?

OpenStudy (anonymous):

yep...and the sequence satisfies the given inequality

OpenStudy (nowhereman):

Then you really should have wrote so in the beginning. Mathematics is all about exactness. And "a sequence" and "example" say something completely different from what you meant...

OpenStudy (anonymous):

i m sorry..i told u.. :) so my proof correct?

OpenStudy (nowhereman):

It still doesn't work out. Let's look at the second element in the sequence. There you have \(n=2\) and \(x_2 = 3\) so for \(m=1\) you have the inequality, but \(m = 1 \neq 2 = [n+0.5]\)